OpenStudy (anonymous):

help!! At a high school, the average grade for the English 10 provincial is 64, with a standard deviation of 10. If 20 students scored between 73 and 86 on the exam, how many students took the exam? Use z scores in your solution

OpenStudy (anonymous):

ok lets solve this \[\Large P(73<X<86)\] \[\Large Z=\frac{x-u}{\sigma}\] so \[\Large (\frac{73-64}{10})<Z<(\frac{86-64}{10})\] so we have \[\Large (0.9<Z<2.2<)\] can you do this using table ?

OpenStudy (anonymous):

@Murry can you find above?

OpenStudy (anonymous):

and it is \[\Large (0.9<z<2.2)\]

OpenStudy (anonymous):

@Murry your ans is 118 ? can you confirm this ?

OpenStudy (anonymous):

How do I find the above? Do i look at the z-score table?

OpenStudy (anonymous):

yes first find for z=2.2 then find z=0.9 subtract the two values

OpenStudy (anonymous):

your answer is 118?

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

did you get it ?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

can you show me your work how did you get it ?

OpenStudy (anonymous):

it shows that the z score should be .9890 ...?

OpenStudy (anonymous):

yes for z=2.2 it is u got it now for z=0.9

OpenStudy (anonymous):

.8389 ?

OpenStudy (anonymous):

little bit different i found z=0.8159

OpenStudy (anonymous):

ok now subtract two values 0.9890-0.8159=0.1731 it means 17.3 % of the 20 are getting no between 73 and 86 so divide 20 by 0.1731 20/0.1731=115.5 =116 students

OpenStudy (anonymous):

Thank you, your a life saver

OpenStudy (anonymous):

you 're welcome :)